how many Cl atoms are in 0.0728 g of PCl3?



1
Expert's answer
2021-09-30T03:24:52-0400

Molar mass of Cl = 35.453

Molar mass of PCl3 = 137.33

= 0.0728/137.33= 0.00053

106.36/137.33×0.00053

= 0.00041m

6.023× 1023×0.00041

= 2.47×1023atoms